# Robust Few-shot Transfer Learning for Knowledge Base Question Answering with Unanswerable Questions

### v1

@@ Abstract
Real-world KBQA applications require models that are (1) robust – e.g., can differentiate between answerable and unanswerable questions, and (2) low-resource – do not require large training data. Towards this goal, we propose the novel task of few-shot transfer for KBQA with unanswerable questions. We present FUn-FuSIC that extends the state-of-the-art (SoTA) few-shot transfer model for answerable-only KBQA to handle unanswerability. It iteratively prompts an LLM to generate logical forms for the question by providing feedback using a diverse suite of syntactic, semantic and execution guided checks, and adapts self-consistency to assess confidence of the LLM to decide answerability. Experiments over newly constructed datasets show that FUn-FuSIC outperforms suitable adaptations of the SoTA model for KBQA with unanswerability, and the SoTA model for answerable-only few-shot-transfer KBQA.
@@ 1 Introduction
The semantic parsing formulation of the Knowledge Base Question Answering (KBQA) task takes a Knowledge Base (KB) and a natural language question as input and requires as output a logical form (or program) that produces an answer upon execution over the KB. KBQA has important real-world applications, which require KBQA systems to be low-resource (requiring only a few task-specific labeled examples) and robust. One important aspect of robustness is the ability to identify unanswerable questions, which arise naturally in real applications Patidar et al. (2023). Traditional supervised models Saxena et al. (2022); Zhang et al. (2022); Mitra et al. (2022); Wang et al. (2022); Das et al. (2022); Ye et al. (2022); Chen et al. (2021); Das et al. (2021); Shu et al. (2022); Gu et al. (2023) and even recent LLM-based few-shot in-context learning (FS-ICL) architectures Li et al. (2023); Nie et al. (2024) are found wanting in both aspects.
Recent work has addressed low-resource KBQA in the form of few-shot transfer learning Patidar et al. (2024) – the idea is to exploit existing KBQA datasets (source) and transfer them few-shot to a target setting. Their FuSIC-KBQA architecture uses supervised KB retrieval followed by generation and refinement of logical forms using LLM few-shot in-context learning.
A parallel direction addresses unanswerable questions for KBQA Patidar et al. (2023). Unanswerability arises due to deficiencies in the schema or the data in the KB.  Faldu et al. (2024) proposes RetinaQA – a supervised architecture (with high training data needs) that is robust to both answerable and unanswerable questions. However, no existing architecture simultaneously addresses the low-resource and robustness requirements.
@@
In response, we propose the task of few-shot transfer learning for KBQA in the unaswerability setting. Specifically, the target KBQA task offers only a few labeled examples, while the source KBQA task has thousands of labeled examples containing answerable questions only.
Building a model for this setting could involve extending FuSIC-KBQA for unanswerability or RetinaQA for few-shot transfer. Since RetinaQA’s design requires large volumes of labeled training data in the target and it fares poorly in few-shot setting, we build our model on top of FuSIC-KBQA. A natural vanilla extension is to modify the LLM prompt for unanswerability and including different types of unanswerable questions in the few shots, as proposed in the broader unanswerable QA literature Slobodkin et al. (2023). Unfortunately, we find that this largely fails for unanswerable questions that do not have a valid logical form – the simple modification regularly labels these as answerable and outputs incorrect logical forms. A second issue is that FuSIC-KBQA prompts LLM to refine the logical form if its execution leads to an empty answer – but this assumes that all questions are answerable and naturally breaks down in the face of unanswerability.
To handle this, we develop FUn-FuSIC (Feedback for Unanswerability in FuSIC-KBQA). FUn-FuSIC takes the vanilla approach further by detecting when the LLM is very uncertain about the logical form, which is suggestive of the question being unanswerable. For this, it employs logical form refinement by providing LLM repeated feedback about potential errors, including syntactic, semantic and execution-related error checks, and a novel strategy of back-translating the logical form to check equivalence with the original question. To assess the model’s lack of certainty about the logical form, FUn-FuSIC adapts self-consistency Wang et al. (2023) – in the absence of enough certainty, schema-level unanswerability is inferred. On the other hand, when a sufficient number of logical forms agree on the answer (empty or non-empty), the LLM is finally prompted to choose the best one, along with inferring answerability (for non-empty answer) or data-related unanswerability (for empty answer).
For evaluation, we first create two new datasets for few-shot transfer for KBQA with unanswerability. Using these, we show that, in aggregate, FUn-FuSIC comprehensively outperforms RetinaQA, when adapted for few-shot transfer, as well as the vanilla extension to FuSIC-KBQA.
We further demonstrate that the broader idea of iterative generation of logical forms, with feedback from a diverse suite of error checks, holds promise for KBQA in general. Using experiments over existing few-shot transfer datasets for answerable KBQA, we find that the restriction of FUn-FuSIC for the answerable setting outperforms FuSIC-KBQA, which is the SoTA KBQA model in this setting.
@@
In summary, our specific contributions are as follows. (a) We propose the problem of few-shot transfer for KBQA with unanswerability. (b) We present FUn-FuSIC that uses iterative feedback for error checks and self-consistency adapted for unaswerability. (c) We create new datasets for the proposed task which we make public  [^1] https://anonymous.4open.science/r/FUn-FuSIC-1704. (d) We show that FUn-FuSIC outperforms adaptation of SoTA KBQA models for our setting. (e) We also show that in the standard answerable setting, FUn-FuSIC outperforms the corresponding SoTA model.
@@ 2 Related Work
There is a lot of work on in-domain KBQA using supervised models  Ye et al. (2022); Chen et al. (2021); Shu et al. (2022); Gu et al. (2023) as well as LLM few-shot approaches Li et al. (2023); Nie et al. (2024); Shu and Yu (2024). These require significant volumes of labeled data to train or select the most appropriate few shots.
@@
Also in the in-domain setting, unanswerability has been studied for KBQA Patidar et al. (2023); Faldu et al. (2024).  Patidar et al. (2023) created the GrailQAbility dataset with different categories of unanswerability, and showed that answerable-only models, when adapted superficially for unaswerability, are inadequate in different ways. RetinaQA Faldu et al. (2024) is the SoTA model for unaswerability. It learns to discriminate between valid logical forms and those rendered invalid due to schema-level incompleteness, and constructs logical forms top-down using sketch-filling to handle data-level incompleteness. However, this also requires large volumes of training data.
In the transfer setting for KBQA, one formulation makes use of large volumes of unlabeled training data in the target Cao et al. (2022); Ravishankar et al. (2022). More recently, FuSIC-KBQA has addressed the few-shot transfer setting with only a small volume of labeled training examples in the target Patidar et al. (2024). This uses a retrieve-then-generate framework using one or more supervised retrievers and LLM few-shot in-context learning based generators. The LLM generation also goes through multiple iterations of execution-guided feedback until a non-empty answer is obtained. However, the transfer setting has only considered answerable questions.
@@
LLM prompting techniques have been used to address unanswerability outside of KBQA Slobodkin et al. (2023), but without feedback for iterative generation. Other approaches  Shinn et al. (2023); Chen et al. (2023b) use execution based refinement for programs but do not involve unanswerability.
In the prompting literature, self-consistency marginalizes over multiple reasoning paths to arrive at a consensus answer Wang et al. (2023); Chen et al. (2023a). This works only when significant probability mass accumulates for at least one answer, which is selected as the winner. This breaks down under unaswerability, when reasoning paths to all non-empty answers have low probability.
3 Problem and Proposed Approach
Figure 1: Illustration of FUn-FuSIC for a specific question when executed over three different KBs. The question is answerable for KB3, but unanswerable for KB1 and KB2, due to schema incompleteness and data incompleteness respectively. The iterations as shown as i=1,2,3,4, The consistency block is below these. Colored boxes show different checks. Additions to the set of candidate logical forms are shown in the gray columns.
We first briefly cover background on KBQA, unaswerability and few-shot KBQA. More details are in the Appendix (Sec. A.1). We then define the few-shot setting with unanswerability and our proposed approach.
3.1 Few-shot Transfer Learning for KBQA with Unanswerability
@@
A Knowledge Base $G$ consists of a schema with data stored under it. The schema consists of entity types $T$ and binary relations $R$ defined over pairs of types. The data consists of entities $E$ as instances of types $T$, and triples or facts $F\subseteq E\times R\times E$. We are given a target Knowledge Base $G^{t}$ and a natural language question $q^{t}$, and the goal is to generate a structured query or a logical form $l^{t}$, which when executed over $G^{t}$ returns an answer set $a^{t}$. We are provided with target few-shots $D^{t}$ containing tens of labeled training examples of questions and logical forms. A related source domain has a source knowledge base $G^{s}$ (with its own types, relations, entities and facts), and a larger source training set $D^{s}$ with thousands of labeled training examples. The source and target KBs and the distributions of questions and logical forms can be very different.
@@
In general, a KB may have both answerable and unanswerable questions Patidar et al. (2023). A question $q$ is answerable for a KB $G$ if it has a corresponding logical form which when executed over $G$ returns a non-empty answer $A$. In contrast, a question is unanswerable for $G$, if it either (a) does not have a valid logical form for $G$ (schema-level unanswerability), or (b) it has a valid logical form for $G$, but which on executing on $G$ returns an empty answer (data-level unanswerability). Schema-level unaswerability arises due to missing types (classes) and relations. On the other hand, data-level unaswerability arises due to missing entities and facts. Absence of any mentioned question entity, however, invalidates the logical form. More details are in the Appendix (Sec.  A.1.3).
@@
We now propose the problem of few-shot transfer learning for KBQA with unanswerability. Here, the target KB $G^{t}$ is incomplete, and target test questions $q^{t}$ may be answerable or unanswerable due to missing schema or data. The few shot examples $D^{t}$ contain both answerable and unanswerable questions of different types. Since almost all large KBQA datasets contain only answerable questions, we assume that the source training data $D^{s}$ contains only answerable questions. So, in addition to the differences between source and target KBQA tasks in the existing transfer learning setting, the unanswerability gap also needs to be bridged.
3.2 Proposed Approach: FUn-FuSIC
Our proposed architecture FUn-FuSIC builds upon and adapts FuSIC-KBQA for unaswerability. We preserve its retrieve-then-generate framework where the first stage is KB retrieval with one or more supervised retrievers, which are source-trained and then target fine-tuned. The retrieval is LLM reranked to increase target relevance. This forms the input to the generative stage, which prompts an LLM to generate the logical form using the retrieval, either in few-shot or zero-shot fashion. More details of these stages are in the Appendix (Sec. F). On top of this FuSIC-KBQA design, FUn-FuSIC makes two modifications. First, the logical form generation instruction includes the possibility of unanswerability, and the few shots include examples of the two types of unanswerable questions. The detailed prompt is in the Appendix (Sec. G.1). Secondly, the generated logical form is refined using feedback as before, but this is now adapted for unaswerability. In the rest of this section, we describe feedback for unaswerability (FUn) in detail.
We observe that the first prong of our solution – prompting for unanswerability (PUn) – largely biases the LLM towards one type of error. For $l=$ NK questions, LLM with PUn mostly generates incorrect logical forms instead of generating NK. The essence of our second prong, FUn, is to assess the LLM’s confidence about any logical form being the correct one for the question. If this confidence is low, then FUn outputs $l=$ NK and $a=$ NA. If there is sufficient confidence in a specific logical form $\hat{l}$, then FUn outputs $l=\hat{l}$. If $\hat{l}$ upon execution returns a non-empty answer $\hat{a}$, then FUn outputs $a=\hat{a}$, else the output is $a=$ NA.
We illustrate flow of FUn using examples in Fig. 1. The complete algorithm is in the Appendix (Sec. 2). FUn assesses confidence in a specific logical form using a suite of checks – syntactic, semantic, and execution-based. Some of these are performed using rules defined for the KB, while others involving uncertainty are performed using LLM prompting. FUn goes over $n$ iterations. In the $i^{th}$ iteration, it generates a new logical form $l^{(i)}$ based on feedback received from checks in earlier iterations. $l^{(i)}$ goes through a sequence of checks. If it passes all checks, FUn outputs $l=l^{(i)}$. If $l^{(i)}$ fails a certain check, it is rejected. On the other hand, if it fails some combination of uncertain checks, it is added to the candidate logical form set $L$. The error message from each failed check is appended to the prompt for the next iteration.
If no clear winner emerges at the end of $n$ iterations, FUn assesses the LLM’s confidence about $L$ as a whole, and also about each logical form in $L$. For this, we propose self consistency with unanswerability (scUn). Let $L_{ne}\subset L$ indicate candidate logical forms that produce non-empty answers upon KB execution. Let $a^{*}$ denote the most popular answer (ties being broken at random) and $L_{ne}(a^{*})\subset L_{ne}$ denote logical forms that produce $a^{*}$. If $|L_{ne}(a^{*})|>t$, where $t$ is a threshold, scUn decides that $a=a^{*}$. LLM prompting then selects the most suitable candidate $l^{*}$ from $L_{ne}(a^{*})$, and FUn outputs ($l=l^{*}$, $a=a^{*}$).
If no consensus emerges from $L_{ne}$, scUn turns to its complement set, $L_{e}=L-L_{ne}$, having empty answers. These represent candidates that agree on $a=$ NA. If $L_{e}$ is not empty, again LLM prompting identifies the most suitable candidate $l^{*}$ from $L_{e}$, and FUn outputs ($l=l^{*}$, $a=$ NA). If on the other hand, $L_{ne}$ is empty, FUn outputs ($l=$ NK, $a=$ NA).
Next, we discuss the different categories of checks. The high level goal is to prune candidate logical forms without knowledge of the gold logical form, the gold answer or even whether the question is unanswerable or not. The specific prompts for the checks and the detailed structures of the feedback containing the error messages are described in the Appendix (Sec. G.2).
(C1) Syntax Error:
The KB engine returns a syntax error upon executing the logical form. This is also a part of EGF in FuSIC-KBQA. Note that this is a certain check – meaning that any logical form that fails this check cannot, with certainty, be the right logical form.
(C2) KB Inconsistency:
The logical form is inconsistent with the KB schema. This can happen in different ways. (C2a) Incompatibility in types: The logical form contains a variable and a connecting relation whose types are incompatible. This error is also checked in FuSIC-KBQA EGF. (C2b) Schema hallucinations: The logical form contains schema elements (types, relations, entities) not in the target KB. (C2c) Type-casting errors: Literals are not correctly type-cast for the target KB, e.g. numeric literals as float for Freebase. All of these are certain checks, and are implemented using rules defined using the logical form and the KB.
(C3) Question Logical form Disagreement:
The generated logical form is syntactically correct and consistent with the target KB but is not semantically equivalent to the original natural language question. This is not a certain check for two reasons. First, different types of ambiguities may exist in the question, ruling out the existence of a semantically equivalent logical form in a specific KB. For example, the question may ask for a person from a country, when the KB only contains the relations born in and works in between these types. Secondly, this check requires a learning based classifier, inevitably having some probability of error.
FUn performs this equivalence check using a novel multi-stage LLM pipeline. (i) The variable names in the logical form are first naturalized, preserving semantics and considering the natural language question, e.g. by replacing ’?x’ with ’?actor’. (ii) The naturalized logical form is back-translated into a natural language question. (iii) The back-translated natural language questions is checked for semantic equivalence with the original natural language question. The first two steps are performed using zero-shot prompting, while the last is performed using few-shots constructed using the target few-shots $D^{t}$. More details about few-shot construction are in the Appendix (Sec. G.3).
(C4) Answer Incompatibility:
If the logical form is syntactically correct and consistent with the target KB, it returns some answer upon execution. This answer may not be compatible with the original question. (C4a) The entity returned as the answer is already contained in the logical form (and therefore mentioned in the question). (C4b) The answer contains at least one entity with an intermediate (compound value) type, and therefore does not correspond to any real-world entity. (C4c) The returned answer is empty. All of these checks can be implemented using rules defined over the logical form, the answer and the KB. Note that while the first two are certain checks, the last is not. In the presence of unaswerability, a question may have $a=$ NA as the gold standard.
We now discuss how FUn identifies candidate logical forms based on these checks and assesses the LLM’s confidence about these.

### v2

# Iterative Repair with Weak Verifiers for Few-shot Transfer in KBQA with Unanswerability

@@ Abstract
Real-world applications of KBQA require models to handle unanswerable questions with a limited volume of in-domain labeled training data. We propose the novel task of few-shot transfer for KBQA with unanswerable questions and contribute two new datasets for performance evaluation. We present FUn-FuSIC – a novel solution for our task that extends FuSIC-KBQA, the state-of-the-art few-shot transfer model for answerable-only KBQA. We first note that FuSIC-KBQA’s iterative repair makes a strong assumption that all questions are answerable. As a remedy, we propose Feedback for Unanswerability (FUn), which uses iterative repair using feedback from a suite of strong and weak verifiers, and an adaptation of self-consistency for unanswerabilty to better assess the answerability of a question. Our experiments show that FUn-FuSIC significantly outperforms suitable adaptations of multiple LLM-based and supervised SoTA models on our task, while establishing a new SoTA for answerable few-shot transfer as well.
Iterative Repair with Weak Verifiers
for Few-shot Transfer in KBQA with Unanswerability
Riya Sawhney, Samrat Yadav, Indrajit Bhattacharya, Mausam Indian Institute of Technology, Delhi, KnowDis AI riya.sawhney@outlook.com, samratya23@gmail.com, indrajitb@gmail.com, mausam@cse.iitd.ac.in
@@ 1 Introduction
The semantic parsing formulation of the Knowledge Base Question Answering (KBQA) task takes as input a Knowledge Base (KB) and a natural language question, and outputs a logical form (or program) that produces the answer upon execution over the KB. KBQA has important real-world applications, which can require systems to be low resource (i.e., trained only with a few task-specific labeled examples), and robust, specifically the ability to identify questions that cannot be answered based on existing KB.
Traditional supervised models (e.g., Ye et al. (2022); Shu et al. (2022); Gu et al. (2023)) and even recent LLM few-shot in-context learning (FS-ICL) architectures Li et al. (2023); Nie et al. (2024) fall short in both aspects. Limited recent work has addressed these independently – in-domain methods for KBQA with unanswerability trained with large labeled data Patidar et al. (2023); Faldu et al. (2024), and FuSIC-KBQA for few-shot transfer assuming questions are answerable Patidar et al. (2024). No existing work simultaneously addresses both desiderata in a single KBQA system.
@@
In response, we propose the novel task of few-shot transfer learning for KBQA with unanswerability. Specifically, the target domain may offer only a few labeled examples of answerable and unanswerable questions, while the source domain has thousands of labeled examples, but containing only answerable questions.
For KBQA transfer, FuSIC-KBQA uses a retrieve-then-generate framework: retrieval of relevant schema and KB snippets followed by an LLM-based generation and a subsequent iterative error guided repair. Specifically, multiple iterations of execution-guided feedback are run until a non-empty answer is obtained – this assumes that all questions are answerable. For our task, we first design FuSIC-KBQA-U, which, to address Unanswerability, eschews the now-inappropriate repair step, and modifies the LLM prompt to accommodate unanswerable questions, along with relevant in-context exemplars. Unlike in studies for unanswerability in general QA Slobodkin et al. (2023), this vanilla adaptation mostly generates incorrect logical forms for unanswerable questions.
As a remedy, we design a novel solution: FUn-FuSIC (Feedback for Unanswerability in FuSIC-KBQA). The key idea is to modify iterative repair, which earlier relied on a single strong verifier for the logical form’s incorrectness, to rely on a suite of strong and weak verifiers, where strong verifiers identify certain errors, whereas weak verifiers identify potential errors in the current logical form.
FUn-FuSIC’s verifiers consider both the logical form and the answer. For answers, non-emptiness check is now a weak verifier, given potentially unanswerable questions. For logical forms, we use strong verifiers to identify obvious syntactic and semantic errors. We also propose a novel verifier: non-equivalence of the original question and the back-translation of the logical form. This is weak, due to potential errors in back-translation as well as in equivalence classification.
Using such iterative weak verification based repair, FUn-FuSIC constructs a set of candidate logical forms. Simply selecting from this set the logical form with majority answer via self-consistency Wang et al. (2023) breaks down in the face of unanswerability. We introduce self-consistency for unaswerability, which additionally checks the likelihood of the majority answer, empty or otherwise, to produce the final output.
For evaluation in our novel task, we create two KBQA transfer datasets with unanswerable questions in the target. Our experiments show that FUn-FuSIC comprehensively outperforms different types of SoTA models suitably adapted for this task, including LLM-based and more traditional models. We further find that iterative repair of logical forms using weak verifiers holds promise for KBQA in general. Using experiments over benchmark datasets for answerable-only few-shot KBQA transfer, we show that the restriction of FUn-FuSIC for the answerable setting improves upon the known SoTA on the task.
@@
In summary, our specific contributions are as follows. (a) We propose the problem of few-shot transfer for KBQA with unanswerability. (b) We present FUn-FuSIC that uses iterative repair with error feedback from a diverse suite of weak verifiers. (c) We create new datasets for the proposed task, which we make public. [^1] https://anonymous.4open.science/r/FUn-FuSIC-1704 (d) We show that FUn-FuSIC outperforms adaptation of SoTA KBQA models for our setting. (e) We also show that for standard answerable-only KBQA, FUn-FuSIC outperforms the corresponding SoTA model.
@@ 2 Related Work
In-domain KBQA using supervised models  Saxena et al. (2022); Zhang et al. (2022); Mitra et al. (2022); Wang et al. (2022); Das et al. (2022); Ye et al. (2022); Chen et al. (2021); Das et al. (2021); Shu et al. (2022); Gu et al. (2023) and using LLM few-shot approaches Li et al. (2023); Nie et al. (2024); Shu and Yu (2024) is well explored in literature. These use high volumes of labeled data, either for training or selecting the most relevant few shot exemplars.
@@
For in-domain KBQA, unanswerability has recently been studied Patidar et al. (2023); Faldu et al. (2024).  Patidar et al. (2023) create the GrailQAbility dataset with different categories of unanswerability, and show the inadequacy of superficial adaptations of answerable-only KBQA models. RetinaQA Faldu et al. (2024) is the SoTA model for KBQA unaswerability. However, this also requires large volumes of training data.
For KBQA transfer Cao et al. (2022); Ravishankar et al. (2022), low-resource was originally not a focus. More recently, few-shot transfer for KBQA has been addressed Patidar et al. (2024). Its FuSIC-KBQA model uses a retrieve-then-generate framework with an LLM-based generation stage with iterative error feedback based repair.
However, this formulation assumes answerability of all questions, as described in Section 1.
@@
Simple LLM prompting techniques have been used to address unanswerability outside of KBQA Slobodkin et al. (2023), but without any notion of feedback or iterative repair. Other approaches  Shinn et al. (2023); Chen et al. (2023b) use execution based refinement for program generation but without any notion of unanswerability.
The extensive recent work on program self-repair using LLMs Olausson et al. (2024) also diverges strongly from our few-shot goal. It assumes availability of unit tests for every test instance. More importantly, as in the answerable KBQA literature, it assumes that a valid program exists for each specification.
To the best of our knowledge, ours is the first work to apply iterative weak repair for KBQA, and to use iterative repair for unanswerability.
3 Background & Problem Definition
Figure 1: Feedback with Unanswerability (FUn) and self consistency for Unanswerability (scUn) for a question $Q$ when executed over different KBs with $\leq 4$ iterations. $Q$ is answerable for KB3, but unanswerable for KB1 (schema incompleteness) and KB2 (data incompleteness). FUn iterations are shown using gray blocks, with Strong Verifiers in red and Weak Verifiers in green. (V1: Syntax Error omitted for brevity) scUn is shown using the blue block.
@@ 3 Background & Problem Definition
A Knowledge Base (KB) $G$ consists of a schema and data. The schema consists of entity types (or classes) $T$ and binary relations $R$ defined over pairs of types. The data consists of entities $E$ as instances of types $T$, and triples or facts $F\subseteq E\times R\times E$. Given a target KB $G^{t}$ and a natural language question $q^{t}$, the goal is to generate a structured query or logical form $l^{t}$ (in a KB query language, such as SPARQL), which when executed over $G^{t}$ returns an answer $A^{t}$ (which is a set in general). Target few-shots $D^{t}$ containing tens of labeled training examples of questions and logical forms. A related source domain has a source KB $G^{s}$ (with its own types, relations, entities and facts), and a larger source training set $D^{s}$ with thousands of labeled training examples.
@@
Following Patidar et al. (2023), a question $q$ is answerable for a KB $G$ if it has a corresponding logical form $l$ which when executed over $G$ returns a non-empty answer $A$. A question is unanswerable, if it either (a) does not have a valid logical form for $G$ (schema-level unanswerability), or (b) it has a valid logical form $l$ for $G$, but $l$ returns an empty answer upon execution on $G$ (data-level unanswerability). Schema-level unanswerability arises due to missing types and relations, while missing entities and facts lead to data-level unanswerability. More details are in Appendix  A.1.3.
@@
We now define few-shot transfer learning for KBQA with unanswerability. A target question $q^{t}$ may be answerable or unanswerable due to missing schema or data in the target KB $G^{t}$. Target few shot examples $D^{t}$ contain both answerable and unanswerable questions of different categories. However, most KBQA datasets contain only answerable questions, we model the source training data $D^{s}$ as containing only answerable questions. Now the unanswerability gap between the source and the target also needs to be bridged. More details are in Appendix A.1.
4 Proposed Approach: FUn-FuSIC
Our proposed architecture FUn-FuSIC adapts FuSIC-KBQA for unaswerability. The high-level algorithm is described in Algo. 1. Preserving the retrieve-then-generate framework, the retrieval stage (line 2) performs KB retrieval for $q^{t}$ using a set $R$ of one or more supervised retrievers. Each retriever $R_{i}$ is source-trained and then target fine-tuned. The retrieval output $r$ of each $R_{i}$ consists of relevant schema elements (types and relations) for $q^{t}$, and data paths emanating from mentioned entities in $q^{t}$. The union of these, along with $q^{t}$, is fed to the generation stage, which uses prompting with an LLM $\mathcal{L}$ to generate logical forms using the target few shots $D^{t}$. More details of the retrieval stage are in the Appendix (Sec. A.6).
Algorithm 1 FUn-FuSIC$(q,G^{t},D^{t},R,V^{s},V^{w},\mathcal{L})$
1:$r=\{\}$
[…]
3:$l=\mbox{PUn}(\mathcal{L},I,q,r,D^{t})$
4:$(e,l,A,L)=\mbox{FUn}(\mathcal{L},q,l,n,V^{s},V^{w},G^{t})$
5:if ($\mbox{{\bf if} }(e)\mbox{ {\bf return}}(l^{*},A^{*})$
6:else $\mbox{{\bf else} {\bf return}}(\mbox{scUn}(q,L,t,\mathcal{L}))$
FUn-FuSIC makes three modifications to the FuSIC-KBQA architecture. First, the LLM generation instruction is modified to admit the possibility of unanswerability, and the few shots are modified to include examples of unanswerable questions. However, this simple approach is error-prone. So, we bias the instruction towards one type of error. Specifically, when uncertain about answerability of the question, Prompting for Unanswerability (PUn) (line 3) instructs $\mathcal{L}$ to generate a (possibly incorrect) logical form instead of $l=\texttt{NK}$. The detailed prompt $I$ is in the Appendix (Sec. A.8).
We now come to the more significant modifications. First, $l^{(0)}$ is iteratively repaired using feedback as before, but this step is adapted for unaswerability. This iterative repair, which we name Feedback for Unaswerability (FUn) (line 4), either confidently outputs a single logical form $l$ (with corresponding answer $A$) (line 5) or generates a set $L$ of candidate logical forms, which is further analyzed. For this, we introduce self-consistency for Unaswerability (scUn). scUn checks the likelihood of the majority answer in $L$, empty or otherwise, to produce the final output. In the rest of this section, we describe FUn and then scUn. Fig. 1 illustrates flow of FUn and scUn using examples. A real example of FUn execution is in Sec. A.13.
Algorithm 2 $\mbox{FUn}(q,l,n,V^{s},V^{w},G,\mathcal{L})$
1:$i=0$, $k=0$, $L=\{\}$, $F=\mbox{""}$"
2:while $++i\leq n$ do
3:     for $j=1\mbox{ to }k_{1}$ do
4:         $(e,f)=V^{s}_{j}(l,q,G)$
5:         $F=\mbox{Append}(F,f)$
6:         $\mbox{{\bf if} }(!e)\mbox{ {\bf break}}$
7:     end for
8:     for $j=1\mbox{ to }k_{2}$ do
9:         $(e,f)=V^{w}_{j}(l,q,G)$
10:         $F=\mbox{Append}(F,f)$
11:         if ($\mbox{{\bf if} }(e)\mbox{ }L=L\bigcup\{l\}$; $k++$
12:     end for
13:     $l=Gen(\mathcal{L},I,q,F)$
14:end while
15:if ($\mbox{{\bf if} }(k=k_{2})\mbox{ {\bf return}}(\mbox{True},l,\mbox{Exec}(l,G),L)$
16:else $\mbox{{\bf else} {\bf return}}(\mbox{False},l,\{\},L)$
The FUn algorithm is described in Algo. 2 Starting with the initial logical form $l^{(0)}$, FUn performs at most $n$ verify-and-repair iterations to create a candidate set $L$ of probable logical forms. The example shows 3 FUn iterations for KB1 and KB2, and 2 iterations for KB3. In the $i^{th}$ iteration, FUn generates a new logical form $l^{(i)}$ by prompting $\mathcal{L}$ using $q^{t}$ and feedback $F$ received from checks in all previous iterations (line 13). $l^{(i)}$ goes through a sequence of verifications. FUn uses two sets of verifiers. The strong verifiers $V^{s}$ are guaranteed to be correct, while the weak verifiers $V^{w}$ are potentially erroneous. The specific verifiers that we use in this paper are defined later in the section. A template-based feedback string $f$ is appended to the generation prompt for $l^{(i+1)}$ based on the specific verifier that $l^{(i)}$ failed. If $l^{(i)}$ fails a strong verifier, it is rejected (line 6). In the example, this happens for all three KBs in iteration 1. If $l^{(i)}$ passes all checks, strong and weak (line 15), FUn terminates by outputting $(l=l^{(i)},A=A^{(i)})$, where $A^{(i)}$ is the answer obtained by executing $l^{(i)}$. This happens in iteration 3 for KB3. Otherwise, if $l^{(i)}$ passes at least one weak verifier but not all, it is added to candidate logical form set $L$ (line 11). This happens for iterations 2, 3 and 4 for KB1 and KB2.
Logical form verifiers:
FUn uses a suite verifiers, categorized as strong ($V^{s}$) and weak ($V^{w}$). These may be syntactic, semantic, or execution-based, defined using simple rules or complex LLM functions over $l$, $q$ and $G$. Note that unlike unit tests in program synthesis, the verifiers do not have knowledge of the gold logical form, the gold answer or answerability of the question.
We now briefly describe the specific verifiers that we use for this paper. Additional details about the verifiers are in the Appendix (Sec. A.8.1 and Sec. A.2). Note however that FUn is a framework that is equally capable of working with a wholly different suite of verifiers.
(V1) Syntax Error:
As in FuSIC-KBQA, this verifier executes the logical form $l$ over $G$ and checks for syntax error. This is a strong check: a valid logical form cannot have syntax error.
(V2) KB Inconsistency:
A logical form $l$ may be inconsistent with the schema of $G$. We identify semantic errors of different categories, such as type-incompatibility and schema hallucinations, implemented using rules over $l$ and $G$. These are also strong verifiers.
(V3) Question - Logical Form Disagreement:
This verifier checks if $l$ is semantically equivalent to the original natural language question $q$. In our example, LF3 for KB1 disagrees with Q. This is a weak verifier. First, $q$ may not have any equivalent logical form for $G$ due to intrinsic ambiguities. For example, $q$ mentions a person from a country, when $G$ has the relations born in and works in between these types. Secondly, this is implemented as a probabilistic classifier that can err. We define equivalence check between $l$ and $q$ using a novel multi-stage LLM pipeline, involving naturalization of $l$ to $l^{n}$, back-translation of $l^{n}$ to natural language question $q^{b}$ and semantic equivalence check between $q$ and $q^{b}$. More details are in Sec.A.2.
(V4) Answer Inconsistency:
This verifier executes $l$ over $G$ to obtain an answer $A$ and then checks its compatibility with $q$. This may fail for different reasons, such as (V4a) $A$ containing an entity mentioned in $q$, (V4b) $A$ being empty, and others. Note that V4a is a strong verifier while V4b is weak, since an empty answer is valid for unanswerable questions (as for LF3 for KB2), but invalid for answerable ones.